Emma is 10 years older than half Olivia"s age If the sum between Emma's age and Olivia;s age is 31 How old is Emma
Answer by Nigelyul(18)

Let Emma's age be "a"
Olivia's age be "b"
a = 10 + (1/2)(b)
a = 10 + b/2
a + b = 31
10 + b/2 + b = 31
3b/2 = 21
3b = 42
b = 14
a = 10 + (1/2)(14)
a = 17
Therefore, Emma's age is 17 years old.